OpenStudy (anonymous):

5y2 – 12y + 7 = (y – 1)( )

OpenStudy (owlcoffee):

Try solving the left part of the equality using the general formula.

OpenStudy (anonymous):

i got y(5y-12)

OpenStudy (owlcoffee):

May I see what you did to obtain taht result?

OpenStudy (mathmate):

When you factorize a quadratic expression ax^2+bx+c=(px+q)(ux+v)=pu x^2+(pv+qu) x +qv By comparing coefficients, we know the following relations: a=pu b=pv+qu c=qv In the given problem, a=5, b=-12 c=5 p=1 q=-1 Find u, v using the relations given above